OpenStudy (anonymous):

differential equations linear equations dy/dx+2xy=x^3 reviewing for test tomorrow. I have problems with two questions of my review please help!

OpenStudy (anonymous):

\[dy/dx+2xy=x^3\] integrating factor is \[e^{(x^2)}\] so...

OpenStudy (anonymous):

\[e^{(x^2)}y=\int\limits e^{(x^2)}x^3dx+c\]

OpenStudy (abb0t):

Wait wait wait a minute, where did you get \(e^{x^2}x^3\)?

OpenStudy (unklerhaukus):

linear integrating factor is \(\large e^{\int2x\mathrm dx}=e^{x^2}\)

OpenStudy (abb0t):

Oh yes yes yes, I see.

OpenStudy (unklerhaukus):

so to evaluate that integral you might want to substitute u=x^2

OpenStudy (anonymous):

yes i tried u sub with u=x^2 but then du=2xdx and we dont have that

zepdrix (zepdrix):

Rearrange things, maybe it'll jump out at you :) \[\Large \int\limits x^2\cdot e^{x^2}(x\;dx)\]

zepdrix (zepdrix):

So the exponent is u. Is there another u somewhere perhaps? :D

OpenStudy (abb0t):

My brain is either warped or damn u gaiz r good :)

zepdrix (zepdrix):

Dat's ok abs +_+ You get prettier every day

OpenStudy (unklerhaukus):

lolz

OpenStudy (abb0t):

:') I knew i made the right choice when I gave out that medal

OpenStudy (anonymous):

ok so u=x^2 du=2xdx 1/2du=xdx so integral becomes.. \[1/2\int\limits u e^{(x^2)}du\]??

zepdrix (zepdrix):

The exponent becomes u also right? :D

OpenStudy (anonymous):

oh yeah

zepdrix (zepdrix):

\[\Large \frac{1}{2}\int\limits u e^{u}\;du\] Mmm good good looks like you're on the right track. That should make integration by parts a bit easier.

OpenStudy (anonymous):

so after tabular its 1/2ue^u-e^u

zepdrix (zepdrix):

Looks good, don't forget to distribute the 1/2.

OpenStudy (anonymous):

\[x^2e^{(x^2)}/2-e^{(x^2)}/2\]

zepdrix (zepdrix):

Good good. We don't want to forget the + C before we do our next step.

OpenStudy (anonymous):

\[y=x^2/2-1/2+c\]

zepdrix (zepdrix):

Woops! That's why we wanted to add the +C before dividing through! XD Don't forget to divide the C. We can't absorb a variable term into the C.

OpenStudy (anonymous):

oh yeah so ce^(-x^2)

OpenStudy (anonymous):

awesome thank you zepdrix now i only have one more problem then i can call it a night!

zepdrix (zepdrix):

cool \c:/

OpenStudy (anonymous):

im about to open my last question if you would like to help real fast....i would appreciate it!